OpenStudy (anonymous):

Solve: | m + 8| = 16 A.{8} B.{-24, 8} C.{8, 16} D.{-24}

OpenStudy (anonymous):

B

OpenStudy (anonymous):

well |m+8| cannot be negative and m=16-8 so m=8 ans = A.

OpenStudy (anonymous):

ummmmm its the absolute value of the whole thing, so if m was -24, it would be |-16| =16

OpenStudy (anonymous):

sure but if m=8 it still works right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so its B

OpenStudy (anonymous):

B) m+8=16 & m+8=-16 m=8&m=-24

OpenStudy (anonymous):

Yes you are right, both are correct and both appear in the set. Answer is B